OpenStudy (toxicsugar22):

One ml of original bacterial culture is put into 199 ml of water in bottle x. What is the final dilution in bottle x?

hartnn (hartnn):

Dilution = 1 ml of sample / (1 ml of sample + x ml of buffer) here, buffer is water with 199 ml so, dilution = 1/ (1+199) = ... ?

OpenStudy (anonymous):

i was never good with biology but this seems pretty simple. It really is just a math equation.

OpenStudy (toxicsugar22):

200 is that it

hartnn (hartnn):

actually, 1/200

OpenStudy (toxicsugar22):

.005

hartnn (hartnn):

yes, 0.005 or 5 * 10^-3

OpenStudy (toxicsugar22):

If 1 ml from bottle x in problem 2 is placed into 199 ml of water in bottle y, what is final dilution in Bottle y

hartnn (hartnn):

i am not sure about this, but i think we just multiply the 2 dilutions to get final dilution, that is, (1/200) * (1/(1+199)) = 1/ (200*200)

OpenStudy (nincompoop):

are we talking about dilution factor?

OpenStudy (nincompoop):

then it is reversed

OpenStudy (nincompoop):

you're doing a concentration factor at the moment

OpenStudy (toxicsugar22):

A student make 1:10000 dilution of bacteria and puts 0.1 ml of this into I petri plate with medium. Te next day, there are 72 colonies growing on the plate. What is the cfu/ml of the original bacterial culture?

OpenStudy (toxicsugar22):

Yes I am niccompoop

OpenStudy (nincompoop):

so let us begin with 1 mL + 99 mL = 100 mL dilution factor = 100 mL / 1 mL the mL cancels out, and you are left with 100. therefore, the dilution factor (DF) is 1:100 dilution (10^2) concentration factor = 1 mL / 100 mL again, the mL cancels out, and you are left with 1/100. therefore, the concentration factor (CF) is 0.01 or 10^-2

OpenStudy (nincompoop):

this was your original problem. I was just pointing out that you were doing concentration factor with the solution you had. the dilution factor is the opposite or reciprocal DF = final volume / aliquot volume where your final volume is aliquot + diluent aliquot is a portion of the original sample. diluent is the material that is used to dilute the sample

OpenStudy (nincompoop):

CF = aliquot volume / final volume

OpenStudy (toxicsugar22):

What about the third question I posted about the students plates

OpenStudy (nincompoop):

okay it says A student make 1:10000 dilution of bacteria and puts 0.1 ml of this into I petri plate with medium. Te next day, there are 72 colonies growing on the plate. What is the cfu/ml of the original bacterial culture? if our dilution is 1:10,000 it must mean that our original sample was 1mL, and that the diluent was 9,999 mL

OpenStudy (nincompoop):

but it is asking you for CF, right?

OpenStudy (nincompoop):

CF = aliquot / final volume

OpenStudy (toxicsugar22):

Yes cfu/ml

OpenStudy (nincompoop):

how would you solve that based on ALL of the information I've given you?

OpenStudy (nincompoop):

if you have answer choices list them and I will explain

OpenStudy (toxicsugar22):

Is the answer 720000

OpenStudy (nincompoop):

how did you get that?

OpenStudy (nincompoop):

I need the solution and logic behind it

OpenStudy (toxicsugar22):

Bacause we were taught cfu/ ml iequal # of colonies times dilution factor

OpenStudy (nincompoop):

is your answer part of choices?

OpenStudy (toxicsugar22):

No

OpenStudy (nincompoop):

colonies * DF = 27 * 10^4 = 2.7 x 10^5 CFU per mL of bacteria

OpenStudy (nincompoop):

my bad, you put 72 … LOL so it is 7.2 x 10^5

OpenStudy (nincompoop):

OMG I went dyslexic for a second, my apologies

OpenStudy (toxicsugar22):

Ok help with one more please

OpenStudy (toxicsugar22):

How much water do you need to make a 1:10 dilutions using 3 ml of your original bacterial culture.

OpenStudy (nincompoop):

but I am doubtful it is straightforward as that. because the amount that was put in the petri dish was 0.1 mL not 1 mL

OpenStudy (toxicsugar22):

But they don't ask for the dilution factor

OpenStudy (nincompoop):

0.1 mL is equivalent to 100,000 or 10^5 dilution factor

OpenStudy (nincompoop):

your dilution factor will dictate your CFU so you need to be careful

OpenStudy (nincompoop):

you essentially have 100,000 DF * 72 = 7,200,000 or 7.2 x 10^6

OpenStudy (toxicsugar22):

Ok what about the last question can you help me with that

OpenStudy (nincompoop):

I don't think you understand it

OpenStudy (nincompoop):

this is why I am hesitant to proceed with the other problem

OpenStudy (toxicsugar22):

I do I got it thanks

OpenStudy (nincompoop):

if you truly understand it, then you should be able to solve the next question

OpenStudy (toxicsugar22):

Is it 1/10 times 3 / 100

OpenStudy (toxicsugar22):

Tat is a start

OpenStudy (nincompoop):

How much water do you need to make a 1:10 dilutions using 3 ml of your original bacterial culture. how about 3:30 = 1:10 see how it works?

OpenStudy (toxicsugar22):

Okay

OpenStudy (nincompoop):

you have 3 mL and you are looking for 1:10 DF the definition of DF again is what?

OpenStudy (toxicsugar22):

Dilution factor

OpenStudy (nincompoop):

well you just gave me what DF stands for… what is the formula?

OpenStudy (toxicsugar22):

1 / dilution and volume

OpenStudy (toxicsugar22):

1/ dv

OpenStudy (nincompoop):

final volume / aliquot = (aliquot + diluent) / aliquot (3 + x) / 3 = 10

OpenStudy (toxicsugar22):

3 plus 30

OpenStudy (nincompoop):

3 + 30 = 33 33/3 ≠ 10

OpenStudy (toxicsugar22):

30

OpenStudy (toxicsugar22):

I put 30

OpenStudy (nincompoop):

(3 + x) / 3 = 10 multiply both sides by 3, you get \[3\times \frac{ 3+x }{ 3 }=10\times3\]

OpenStudy (toxicsugar22):

27 plus 3

OpenStudy (nincompoop):

there you go

OpenStudy (nincompoop):

that means you would add 27 mL of diluent to a 3 mL of aliquot of the sample in order to make a 1:10 dilution

OpenStudy (toxicsugar22):

Mthanks

OpenStudy (toxicsugar22):

Fan me

OpenStudy (nincompoop):

i hope I won't be seeing you ask the same type of questions again :P

OpenStudy (nincompoop):

http://inst.bact.wisc.edu/inst/index.php?module=book&type=user&func=displayarticle&art_id=103 you will use dilution and concentration in both biology and chemistry so it is important you learn it well.